A sample of final exam scores is normally distributed with a mean equal to 22 and a variance equal to 25.

  • Part (a)

    What percentage of scores are between 17 and 27? (Round your answer to two decimal places.)
    %

  • Part (b)

    What raw score is the cutoff for the top 10% of scores? (Round your answer to one decimal place.)

  • Part (c)

    What is the proportion below 14? (Round your answer to four decimal places.)

  • Part (d)

    What is the probability of a score less than 30? (Round your answer to four decimal places.)
Answer #1

Solution :

Given that ,

mean = \mu = 22

standard deviation = \sigma = 5

Part (a)

P(17 < x < 27) = P[(17 - 22)/ 5) < (x - \mu ) /\sigma  < (27 - 22) / 5) ]

= P(-1 < z < 1)

= P(z < 1) - P(z < -1)

= 0.8413 - 0.1587

= 0.6826

percentage = 68.26%

Part (b)

Using standard normal table ,

P(Z > z) = 10%

1 - P(Z < z) = 0.1

P(Z < z) = 1 - 0.1

P(Z < 1.28) = 0.9

z = 1.28

Using z-score formula,

x = z * \sigma + \mu

x = 1.28 * 5 + 22 = 28.4

Raw score = 28.4

Part (c)

P(x < 14) = P[(x - \mu ) / \sigma < (14 - 22) / 5]

= P(z < -1.6)

= 0.0548

proportion = 0.0548

part (d)

P(x < 30) = P[(x - \mu ) / \sigma < (30 - 22) / 5]

= P(z < 1.6)

= 0.9452

probability = 0.9452